A photon travels out of a material with an index of refraction 1.6 into one with an index of refraction 2.4.

(a) What is the speed of light in the first material?

(b) What is the speed of light in the second material?

(c) Say a box moving at 3.00 m/s were to slide up a ramp, and when it levels off its speed has reduced as much as the speed of the photon reduces from part (a) to part (b). How high would the ramp be?
Answer #1

m, = 1.6 8 m2 = 2.4 ne, af. = 3 X 108 ms -1.875 108 m/s 1.6 V = = 3X108 m/s = 1.25 X108 m/s 2.4 - Here vi = 3.0m/s mp = funkt